# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v2

@@ III.4 Model of Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [Qian2022], supplemented with insights from [Meirzada2018]. This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0. Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states. The model considers 8 states as illustrated in the schematic energy level diagram in Fig. 8.
The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [Qian2022]. The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following [Meirzada2018]. Values used in this work for the rates are provided in the appendix, Table 3.
The model is based on the following assumptions:
  • Steady-State Approximation: The populations reach a steady state within the simulation time, allowing us to analyze the equilibrium behavior.
  • Neglect of Higher-Order Processes: Multi-photon processes and higher excited states are neglected due to their low probabilities under the experimental conditions.
  • Constant Rate Coefficients: The rate coefficients are assumed to be constant and independent of the local temperature and environmental factors.
  • No External Fields: Effects of external magnetic or electric fields on the transition rates are not included.
@@
Figure 8: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), quartet states (P5 and P6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
The simulation results reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases. This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-. Specifically, up to an CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination.
However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
The total fluorescence intensity from the NV centers decreases with increasing CW NIR power (Fig. 9). This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers. The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant and we observe a saturation in the fluorescence.
The $\sigma_{0-}$ ratio increases with CW NIR power. At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig. 9) and $\sigma_{0-}$ ratio (Fig. 6a,c)) observed experimentally.
Figure 9: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.)
III.5 Temperature Simulation
To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air).
@@ III.5 Temperature Simulation
To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond (radius $R=35\;\text{nm}$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
@@
The simulation is based on the following assumptions:
  • Uniform Temperature Distribution: The FND is assumed to have a uniform temperature at any given time due to its small size.
  • Negligible Radiation Losses: Heat loss due to thermal radiation is neglected because it is insignificant compared to conduction at the temperatures considered.
  • Constant Thermal Properties: The thermal properties of the FND and the medium are assumed constant over the temperature range studied.
  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient is considered constant and based on the thermal conductivity of the medium.
  • No Phase Change: The medium remains in the same phase (liquid for water, gas for air) throughout the simulation.
@@
In Water: The temperature increase of the FND is minimal due to the high thermal conductivity and heat capacity of water. The simulation indicates that during the 500 ns NIR pulse:
  • The maximum temperature rise ($\Delta T_{\text{max}}$) is approximately 2 K above the ambient temperature.
  • The temperature increases steadily during the laser pulse and rapidly returns to ambient temperature within nanoseconds after the laser is turned off.
In Air: The temperature increase is more pronounced because air has lower thermal conductivity and heat capacity. The simulation results show:
  • The maximum temperature rise reaches approximately 41 K above the ambient temperature during the 500 ns NIR pulse.
  • The temperature increases during the laser pulse and decreases more slowly than in water after the laser is turned off.
The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments, cf. Fig. 7. In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for. These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
A direct comparison of Fig. 6c)–d) with Fig. 6a)–b) shows that the environment (air versus water) is the controlling parameter. In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively. By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig. 6a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat removal.
For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\lesssim 10^{-4}$ nm) [Acosta2010, Rogers2015]. Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e. $\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold. Values for the measured absorption cross-sections decrease from $\sigma\!\approx\!2\times 10^{-19}$ cm2 on resonance to below $10^{-21}$ cm2 at this detuning [Acosta2010], so direct optical pumping of the singlet transition by the 1064 nm light is negligible. The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
IV Sensing Capabilities of Trapped FNDs
As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature. Measurements included ODMR, fluorescence relaxation, and optical spectrum analysis under variation of the three parameters mentioned above. Results are summarized in Fig. 10.
@@ IV Sensing Capabilities of Trapped FNDs
IV.1 pH Sensing
@@
  • Fluorescence relaxation: A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 10.h). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [Fujisaku2019, Rendler2017].
  • Optical Spectrum: A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig. 10.e and Fig. 11) suggests that alkaline conditions favor the NV- charge state. It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface. Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state [karaveli2016]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig. 11) shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant changes in resonance frequency (Fig. 10.b) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
The $\sigma_{0-}$ ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
To further explore the dependence of the NV charge state on pH, Fig. 11 presents the evolution of the $\sigma_{0-}$ ratio when varying pH from 13 to 1 for the same FND, both under trapping conditions and without NIR laser exposure. The trapped FNDs in suspension exhibit a significantly stronger modulation of $\sigma_{0-}$ compared to the same particle measured on glass. This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing. The pronounced increase in NV center fraction under alkaline conditions (lower $\sigma_{0-}$) confirms the surface deprotonation hypothesis, while acidic conditions suppress this stabilization. Overall, these results validate that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
IV.2 Paramagnetic Species Sensing
@@
  • Fluorescence relaxation: A significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig. 10.g). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [Martinez2020, Rendler2017].
@@
  • Optical Spectrum: A decrease in $\sigma_{0-}$ was observed (Fig. 10.d and Fig. 12). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [Rondin2010]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig. 12) shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant shifts or changes in resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig. 10.a).
@@
These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+. The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding. The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from fluorescence relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
IV.3 Temperature Sensing
The temperature response of NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
  • ODMR: A temperature-dependent shift in the resonance frequency was observed, with lower frequencies at higher temperatures (Fig. 10.c). This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ [Acosta2010].
  • Fluorescence relaxation: No significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 10.i).
  • Optical Spectrum: No significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 10.f).
ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
V Discussion

### v3

@@ 2.5 Simulations - Model of Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model Qian et al. (2022), supplemented with insights from Ref. Meirzada et al. (2018). This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation. The model considers eight states as illustrated in the schematic energy level diagram in Fig. 2. These states include the negatively charged NV- and the neutrally charged NV0, with each their ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states. The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state $i=1,\ldots 8$ Qian et al. (2022). The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following Ref. Meirzada et al. (2018). Values used in this work for the rates are provided in the appendix, Table 3, and rate-equations are explicitly written in the Supplementary information. Simulations were carried out using the solve_ivp function from the sciPy library in python.
The model assumes that populations reach a steady state within the simulation time, and that transitions happen with constant rate coefficients, independent of local temperature and other environmental factors including external magnetic or electric fields. Further, the model neglects multi-photon processes as well as higher excited states as the probabilities for these are negligible under the experimental conditions.
@@
Figure 2: Energy level diagram of the NV center, showing the ground states $^{3}A_{2},m_{s}=0$ (denoted as 1), $^{3}A_{2},m_{s}=\pm 1$ (2) and $^{2}E$ (7), excited states $^{3}E,m_{s}=0$ (3), $^{3}E,m_{s}=\pm 1$ (4) and $^{2}A$ (8), singlet shelving states $^{1}E$ (5) and $^{1}A_{1}$ (6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
2.6 Simulation - Thermal response
@@
To model a potential temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air). To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond of radius $R$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
@@
Because of the small size of the FNDs, the simulation assumes a uniform temperature throughout the FND at any given time, it assumes insignificant heat loss due to thermal radiation, and constant thermal properties of the FND and the medium, including a constant convective heat transfer coefficient based on the thermal conductivity of the medium. Further, the model assumes that the medium remains in the same phase (liquid) throughout the simulation.
3 Results
Our initial characterisation experiments encompass ODMR and photoluminescence (PL) spectroscopy, and fluorescence relaxometry, with and without NIR irradiation. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation while relaxometry experiments applied the pulse protocol in Fig. 1d). Three conditions were investigated, either with FNDs on glass under dry conditions or with a drop of water, or FNDs in aqueous solution in the optical trap, cf. Fig. 1a)-c). The dry conditions facilitates the understanding of thermal phenomena, such as localized heating effects, and are intended as a clear reference to compare with conditions involving FNDs suspended in aqueous environments. We acknowledge that the dry condition itself lacks direct practical application in sensing contexts. For each set of experiments, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW, evaluated at the position of the sample. Fig. 3 show five representative examples of variation in properties from individual FNDs as NIR power increases. As detailed below, data for FNDs on glass with a drop of water are not shown, but will be described.
Figure 3: Representative examples of variation for individual nanodiamonds. (a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
Fig. 4 summarizes the results for 30 FNDs under dry conditions or in aqueous solution in the optical trap. Examples of the underlying measurements are provided in the Supplementary information.
Figure 4: Summary statistics. a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 1d) for fluorescence relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
To evaluate thermal effects, ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz. Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power. For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW. We also observe asymmetries in ODMR contrasts corresponding to spin transitions $\Delta m_{s}=\pm 1$, in correspondence with earlier results in literature for diamonds with ensembles of NV-centers Mittiga et al. (2018); Broadway et al. (2019); Bauch et al. (2020). Guided by these insights from literature, we interpret the asymmetries to be the result of local strain heterogeneity and uneven charge distributions near the diamond surface.
To evaluate charge-state balance, we record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power. Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity. The curves for FNDs on glass remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30 mW to 60 mW and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water. To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively. Compared to the analysis used in Ref. Sow et al. (2020), we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$. In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR. This is illustrated by the representative examples of fitted values for $T_{\rm relax}$ in Fig. 3b) and the summary of the same for 30 individual FNDs in Fig. 4 parts a) and c).
Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$. For FNDs dried on glass substrates, we observed a significant decrease in the $T_{\rm relax}$ relaxation times with increasing NIR laser power. For FNDs trapped in water, the decrease in $T_{\rm relax}$ due to NIR irradiation was significantly less, and we hardly observe any visible change in the two relaxometry curves. Representative examples are shown in Fig. 3c) with statistics of all results summarized in Fig. 4, part b) and d).
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. For FNDs on glass, dry or with water, three groups were compared (NIR power of 0 mW, 30 mW or 60 mW) while for FNDs in the optical trap, only two groups (NIR power of 30 mW and 60 mW). The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
p-value
FNDs dried on glass
$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
ODMR Frequency Shift <0.0001<0.0001
$\sigma_{0-}$ Ratio 0.0017
FNDs on glass with water
$T_{\rm relax}$ Relaxation Time $0.020$
ODMR Frequency Shift $0.34$
$\sigma_{0-}$ Ratio $0.019$
FNDs Trapped in Water
$T_{\rm relax}$ Relaxation Time 0.221
ODMR Frequency Shift 0.073
$\sigma_{0-}$ Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air. For FNDs on glass substrates after adding water, the p-values of $\simeq 0.02$ for $T_{\rm relax}$ and $\sigma_{0-}$ indicate statistical significance while the ODMR frequency shift, indicative of thermal effects, does not show statistically significant variation. In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), for all three quantities, suggesting that the differences observed across different NIR power levels are not statistically significant.
Overall, as showed in Fig. 3 and Fig. 4, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 1d) and Ref. Russell et al. (2018), potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
The rate equation simulation reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases. This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-. Specifically, up to a CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination. However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0. Consequently, the total fluorescence intensity from the NV centers first increases then decreases with increasing CW NIR power (Fig. 5). This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers. The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant. At low CW NIR powers, the $\sigma_{0-}$ ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence. The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig. 5) and $\sigma_{0-}$ ratio (Fig. 4a,c)) observed experimentally.
Figure 5: Total fluorescence intensity from NV centers as a function of NIR laser power. Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.
@@ 3 Results
The results of the simulation of thermal effects align with experimental observations where NIR illumination causes notable thermal effects in air and much less in aqueous environments, cf. Fig. 3-Fig. 4. A direct comparison of Fig. 4c)–d) with Fig. 4a)–b) shows that the environment (air versus water) is the controlling parameter when it comes to the thermal response. In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively. By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig. 4a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat exchange.
For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\sim 10^{-4}$ nm) Rogers et al. (2015). Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e. $\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold. The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore most likely arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
3.1 Sensing Capabilities of Trapped FNDs
In addition to the multi-modal characterization presented above, the potential of optically trapped FNDs with NV centers for sensing applications were addressed. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature, and the same three measurement modalities (PL spectroscopy, ODMR, and optical relaxometry) as before. Results are summarized in Fig. 6.
@@ 3.1 Sensing Capabilities of Trapped FNDs
The temperature response of nanodiamonds with NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K. As expected, a temperature-dependent shift in the ODMR resonance frequency was observed, with lower frequencies at higher temperatures (Fig. 6c)). This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ Acosta et al. (2010). On the contrary, no significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 6i)). Further, no significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by temperature within the range investigated here (Fig. 6f)). ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements.
@@
No significant changes in ODMR resonance frequency (Fig. 6b)) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states. A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 6h)). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics. A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig. 6e)) and Fig. 7) suggests that alkaline conditions favor the NV- charge state. It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface. Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state Karaveli et al. (2016). The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig. 7) shows a clear advantage of using FNDs in suspension for sensing. Another advantage of the optical trap is that effects of individual variation between FNDs from a batch is reduced since measurements can be made relative to an initial measurement with the same FND.
To further explore the dependence of the NV charge state on pH, Fig. 7 presents the evolution of the $\sigma_{0-}$ ratio when varying pH from 13 to 1 for the same FND, both under trapping conditions and without NIR laser exposure. The trapped FNDs in suspension exhibit a significantly stronger modulation of $\sigma_{0-}$ compared to the same particle measured on glass. This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing. Overall, the results confirm that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
@@
No significant shifts or changes in ODMR resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig. 6a)). A decrease in $\sigma_{0-}$ was observed (Fig. 6d) and Fig. 8). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 Rondin et al. (2010). The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig. 8) shows a clear advantage of using FNDs in suspension for sensing. Finally, a significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig. 6g)). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment Tetienne et al. (2013); Rendler et al. (2017); Martinez et al. (2020).
@@
These findings demonstrate the potential of optically trapped nanodiamonds with NV centers for detecting paramagnetic species like Gd3+. The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from fluorescence relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
4 Discussion